# Cascading Large Language Models for Salient Event Graph Generation

### v1

@@ Abstract
Generating event graphs from long documents is challenging due to the inherent complexity of multiple tasks involved such as detecting events, identifying their relationships, and reconciling unstructured input with structured graphs. Recent studies typically consider all events with equal importance, failing to distinguish salient events crucial for understanding narratives. This paper presents CALLMSAE, a CAscading Large Language Model framework for SAlient Event graph generation, which leverages the capabilities of LLMs and eliminates the need for costly human annotations. We first identify salient events by prompting LLMs to generate summaries, from which salient events are identified. Next, we develop an iterative code refinement prompting strategy to generate event relation graphs, removing hallucinated relations and recovering missing edges. Fine-tuning contextualised graph generation models on the LLM-generated graphs outperforms the models trained on CAEVO-generated data. Experimental results on a human-annotated test set show that the proposed method generates salient and more accurate graphs, outperforming competitive baselines. 111For access to experimental code and data, please refer to: https://github.com/Xingwei-Warwick/CALLMSAE
@@ 1 Introduction
Recent studies on contextualised event graph generation have focused on fine-tuning language models to generate linearised graphs from documents in an end-to-end manner Madaan and Yang (2021); Tan et al. (2024). These methods rely on distant supervision, such as events and event temporal relations detected using an approach called CAEVO McDowell et al. (2017), due to the data-intensive nature of language models and heavy manual efforts of annotating event graphs. However, CAEVO has limitations. It typically considers predicates (e.g., verbs) in text as events and tends to extract many insignificant events, such as “say” and “think”, which add little value to narrative understanding and have minimal connections to other events, thus introducing noise to the event graphs.
To improve the quality of distant supervision graphs, it is essential to consider the saliency of events. We found that CAEVO-extracted events often have low saliency because CAEVO takes a bottom-up approach to event extraction, classifying each predicate as an event or not. In contrast, identifying salient events requires a top-down approach. Existing studies on identifying salient events or entities use the summarisation test to guide human annotation, where an event or entity is considered salient if a human-written summary is likely to include it Dunietz and Gillick (2014); Liu et al. (2018). Given that instruction fine-tuned LLMs perform on par with human writers in news summarisation Zhang et al. (2024), we propose generating salient events by instructing LLMs to first summarise documents before identifying salient events.
@@
We propose CALLMSAE, a CAscading Large Language Model framework for SAlient Event graph generation, serving as a distant signal generator for contextualised graph generation models. We also propose a novel contextualised evaluation metric for comparing salient event graphs.
We provide a large-scale LLM-generated salient event graph dataset ( documents) with three relation types for distant supervision, along with a human-annotated test set ( documents).
We present an extensive experimental evaluation on LLM-generated event relation graphs in terms of event saliency and event relation on the NYT corpus, demonstrating how higher quality salient event graphs can improve contextualised graph generation.
@@ 2 Related Work
The early idea of event relation graph construction comes from UzZaman et al. (2013), which introduces a dataset for evaluating an end-to-end system which takes raw text as input and output TimeML annotations (i.e., temporal relations). CAEVO McDowell et al. (2017) and Cogcomptime Ning et al. (2018) both utilise a wide range of manually designed features to train MaxEnt and averaged perception for extracting events and relations. Han et al. (2019b) proposed a joint event and relation extraction model based on BERT Devlin et al. (2019) and BiLSTM Panchendrarajan and Amaresan (2018). Other researchers focus on developing specialised sub-systems to classify extracted event pairs for relations Ning et al. (2019); Han et al. (2019a); Wang et al. (2020); Tan et al. (2021). ATOMIC Sap et al. (2019) is a large-scale commonsense knowledge graph containing the causes and effects of events. MAVEN-ERE Wang et al. (2022) is built with event coreference, temporal, causal and subevent relations. However, ATOMIC and MAVEN-ERE completely rely on crowdsourcing and thus are difficult to extend. MAVEN-ERE is less than half the size of our dataset and does not consider the saliency of events.
@@ 3 Cascading LLMs to Generate Salient Event Graphs
CALLMSAE combines various prompts in a pipelined manner to generate salient event graphs. In this section, we will first introduce the prompts for generating salient events. Then, we will describe the method for generating relation graphs based on the salient events. Lastly, we define an evaluation metric for comparing event graphs: Hungarian Graph Similarity.
@@ 3.1 Generate Salient Events
The summarisation test (as mentioned in Section 1) is often used to guide the annotation of salient events or entities Dunietz and Gillick (2014); Liu et al. (2018). These studies identify events or entities included in human-written summaries as salient. Similarly, we instruct LLMs to generate a summary first and then extract events from it. Examples of the summarisation and salient event generation prompts are shown in the Appendix (Table 9 and 10).
@@ 3.2 Generate Graphs as Code Completion
While LLMs can extract salient events, they often struggle with identifying event relations Chan et al. (2023); Tan et al. (2024). Prompt engineering for extracting event relations is complex due to the need to incorporate various terminologies and graph constraints. Moreover, prompt efficiency is crucial as generating a large-scale dataset with LLMs can still incur significant computational costs, albeit less than crowdsourcing.
In our method, the main prompt for generating the event relation graph is formulated as a Python code completion task. The graph is defined using the NetworkX222https://networkx.org/documentation/stable/ package in Python, with nodes representing the salient events generated in Section 3.1. LLMs are instructed to complete the code by adding relation edges using NetworkX’s APIs.
Recent research suggests that formulating prompts as code can enhance LLMs’ reasoning abilities Wang et al. (2023); Zhang et al. (2023). In our task, the Python code format effectively incorporates all necessary terminologies, enabling LLMs to understand them without confusion. The Python code format also allows for the inclusion of constraints (e.g., ensuring the graph is a directed acyclic graph) and additional instructions (e.g., ask for explanations) as comments. LLMs can generate explanations as comments without disrupting the main content of the graph. Moreover, the code template simplifies parsing the response, as LLMs are directed to use the “.add_edge()” function to add the relations.
@@ 3.4 Complement Relation Types
Hierarchical, temporal, and causal relations are not independent of each other. We found that if one type of relation depends on another, providing the graph for the first relation can benefit the generation of the dependent relation’s graph. Specifically, we predict the hierarchical relation graph first. Then, we provide this graph to the LLMs and ask them to generate the temporal relation graph. Lastly, with both the hierarchical and temporal relation graphs available, the LLMs predict the causal relation graph.
The hierarchical relation describes two closely related events at different granularity levels. It focuses on the inherent semantics of the events and does not depend on other relation types. For example, “writing a dissertation” is a subevent of “doing a PhD”. Therefore, we choose to predict the hierarchical relations first.
Temporal relations can depend on hierarchical relations. For example, knowing “doing a PhD” happened before “being prompted to Professor” allows us to deduce that “writing a thesis” also happened before “being prompted to Professor”. Thus, we predict temporal relations after hierarchical relations.
Lastly, causal relations depend on both hierarchical and temporal relation, as the antecedent event in a causal relation must occur before the consequence. Therefore, the causal relation is predicted in the last step. For more details about the entire prompting process, please refer to the descriptions and pseudocode in Appendix C.
@@ 3.5 Hungarian Graph Similarity
It is challenging to compare event relation graphs generated by LLMs due to the abstractive nature of generation, making it difficult to align the generated events with the gold standard events Li et al. (2023). Moreover, salient events are often high-level and abstract rather than fine-grained and concrete, which means some variations in wording are not only acceptable but also expected. Instead of using exact matching Zhao et al. (2024) or rule-based token matching Tan et al. (2024) on events and relations to calculate , adopting semantic-based evaluation metrics is more reasonable and fair. As more tasks adopt text generation frameworks, many researchers are also turning to metrics based on language models rather than traditional token matching metrics like ROUGE and BLUE Goyal et al. (2022); Pratapa et al. (2023).
In this study, we propose a novel metric for evaluating LLM-generated event graphs, called Hungarian Graph Similarity (HGS). The metric is based on the Hungarian assignment algorithm Kuhn (1955), which is widely used in the object detection to match generated objects and target objects Carion et al. (2020). It can find the optimal assignment given a cost matrix containing the distance between elements in two lists of objects. We adapt this algorithm to match predicted edges with edges in the gold standard graphs as follows:
@@ 4 Dataset
In this section, we describe how we construct the distant supervision dataset and a human-annotated dataset from the New York Times (NYT) corpus.
@@ 4.2 Annotation Settings
We recruited annotators from Prolific333http://www.prolific.com. There are two subtasks: salient event identification and event relation identification. In the first subtask, the participants are asked to identify the salient event triplets: actor, trigger, and object (optional). We provide the definition of an event and several examples in the guidelines. They are instructed to do the summarisation test: the salient events should be the events they would include in the summary of the given article. Moreover, we provide some prominent features for helping annotators to identify salient events Choubey et al. (2018); Jindal et al. (2020):
@@ 5 Experiments
5.1 Model Settings
We compare our proposed approach with the following baselines:
CAEVO McDowell et al. (2017) is a pipeline system based on a Maximum Entropy (MaxEnt) classifier and manually designed features for extracting events and temporal relations.
Madaan and Yang (2021) trained language models on CAEVO-generated linearised graphs with the language modelling objective. We implemented their method to train a Flan-T5 model.
Han et al. (2019b) proposed a joint event and temporal relation extraction model. We adapted the model to predict hierarchical and causal relations by training it on the MAVEN-ERE dataset Wang et al. (2022). We also replaced BERT with Longformer Beltagy et al. (2020) to enable it to process long documents.
GPT-4 OpenAI et al. (2024) and GPT-3.5444https://platform.openai.com/docs/models/gpt-3-5-turbo are based on the generative pre-train framework. We used “gpt-4-1106-preview” and “gpt-3.5-turbo” respectively.
Mixtral is an LLM based on the Mistral model and the mixture of expert framework. We used the Mixtral 8x7B instruct version Jiang et al. (2024).
Llama3 is an LLM based on the Llama framework555https://ai.meta.com/blog/meta-llama-3/. We used the Llama3-70B-instruct 8-bit version provided by ollama666https://ollama.com/library/llama3:70b-instruct-q8_0. The 8-bit quantization is shown to be degradation-free Dettmers et al. (2022).
We fine-tuned a Flan-T5-base (M) with the relation graphs generated by CALLMSAE, following the same method as in Tan et al. (2024). The baseline prompt evaluates whether each event pair is supported by the document, akin to the hallucination grader described in Section 3.3. Thus, it serves as an ablation of our method without incorporating the code prompt.
5.2 Event Saliency Evaluation
| Mean event | Event | First | Stretch | |
| number | frequency | appearance | Size | |
@@ 5.2 Event Saliency Evaluation
| Code Prompt (dependent rels) | |||||||||
@@ 5.3 Salient Event Relation Graph Evaluation
The salient event relation graph evaluation results are shown in Table 5. Even with the most basic prompting (Baseline Prompt), which queries the relation of each event pair, Llama3 outperforms all the baseline methods on all relation types. However, Baseline Prompt is slow and costly because the number of prompts it needs for building one graph is , where is the number of events in the document. On the other hand, Code Prompt only needs . Moreover, Code Prompt’s overall HGS is significantly higher than Baseline Prompt on all relation types. Baseline Prompt check the event pairs more thoroughly and thus have higher recall but its precision is much lower. The complete CALLMSAE combines the code prompt and hallucination grader for iterative refinement, checking missing relations and verifying them to prevent hallucination. It significantly increases the precision and strikes a balance with recall.
@@ 5.4 Format Error and Cycles in the Graphs
A format error occurs when the generated code blocks fail to pass the Python interpreter. We detected these errors by executing the generated code. If the Python interpreter returns an error, it is classified as a format error. We specified the relation graphs as directed acyclic graphs in the prompt. If there is a cycle in the generated graph, it means that the LLM failed to follow the instructions. A cycle also indicates a violation of logic constraints because all the relations in the event relation graphs are asymmetric. We detected the cycles using the find_cycle() from the NetworkX after obtaining the transitive closure of the graphs.
We prompt each LLM three times on the annotated test set. Table 6 shows the average number of documents encountering format errors or cycles. All LLMs have low rates of format errors which shows that state-of-the-art LLMs can understand the instruction well and generate executable Python code. Among them, GPT-3.5 and Llama3 achieve zero errors. The occurrence of cycles can serve as an indicator of the reasoning ability of the LLMs. About of graphs generated by GPT-3.5 have cycles, suggesting that GPT-3.5 may have limited reasoning ability compared to other LLMs. GPT-4 and Mixtral both have low rates of cycle occurrence, but they are beaten by Llama3 which has no cycle in all generations, showing its remarkable understanding of the transitive and asymmetric constraints in the complex event relation graphs.
| Format Error | Cycle | |
| GPT-3.5 | ||
| GPT-4 | ||
| Mixtral | ||
| Llama3 |

### v2

@@ Abstract
Generating event graphs from long documents is challenging due to the inherent complexity of multiple tasks involved such as detecting events, identifying their relationships, and reconciling unstructured input with structured graphs. Recent studies typically consider all events with equal importance, failing to distinguish salient events crucial for understanding narratives. This paper presents CALLMSAE, a CAscading Large Language Model framework for SAlient Event graph generation, which leverages the capabilities of LLMs and eliminates the need for costly human annotations. We first identify salient events by prompting LLMs to generate summaries, from which salient events are identified. Next, we develop an iterative code refinement prompting strategy to generate event relation graphs, removing hallucinated relations and recovering missing edges. Powered by CALLMSAE, we present NYT-SEG, a large-scale automatically annotated event graph dataset which can serve as distant supervision signals. Fine-tuning contextualised graph generation models on NYT-SEG outperforms the models trained on CAEVO data. Results on a human-annotated test set show that the proposed method generates salient and more accurate graphs, outperforming competitive baselines. 111Code and data: https://github.com/Xingwei-Tan/CALLMSAE.
@@ 1 Introduction
Existing studies on contextualized event graph generation primarily focus on fine-tuning language models for end-to-end generation of linearized graphs from documents Madaan and Yang (2021); Tan et al. (2024). These methods rely on distant supervision, such as events and event temporal relations detected using CAEVO McDowell et al. (2017), due to the data-intensive nature of language models and the significant manual effort required for annotating event graphs. However, CAEVO’s bottom-up, predicate-centric extraction often results in sparse, low-quality event graphs populated with non-salient events. For example, CAEVO frequently identifies all verbs as events, including trivial ones like "say" and "think," which have minimal connections to other events and contribute little to narrative understanding, thus introducing noise. In recent research involving human annotation Dunietz and Gillick (2014); Liu et al. (2018), a summarisation test is leveraged to identify salient events or entities, where an event or entity is considered salient if a human-written summary is likely to include it.
Motivated by these insights, we aim to improve the quality of distant supervision and mitigate noise in event graph generation by incorporating event saliency. Our hypothesis is that effective graph generation benefits from a top-down strategy, where the main events are identified first, rather than extracted solely in a bottom-up manner based on the predicate. Capitalizing on the powerful summarization capabilities of large language models (LLMs), we propose a method that first instructs LLMs to summarize documents before identifying salient events. Specifically, we introduce a code-prompting strategy for constructing event relation graphs encompassing hierarchical, temporal, and causal relations222We extend beyond the CAEVO’s temporal-only relations to encompass multiple relation types. (see Figure 1). Unlike vanilla prompting, which queries each potential event pair individually, our code prompt format generates each relation type in a single pass. Furthermore, we incorporate an iterative refinement process using a hallucination grader to filter spurious edges and iterative generation to recover missing ones. Finally, the abstractive nature of salient events presents an evaluation challenge, as they rarely match gold standard events exactly, even when semantically equivalent. To address this, we propose evaluation metrics based on semantic text embeddings for assessing the generated event relation graphs.
@@
We propose CALLMSAE, a CAscading Large Language Model framework for SAlient Event graph generation, serving as a distant signal generator for contextualised graph generation models.
We propose a novel contextualised evaluation metric for comparing salient event graphs. Our extensive experimental evaluation on LLM-generated event relation graphs in terms of event saliency and event relation on the NYT corpus, demonstrating how higher quality salient event graphs can improve contextualised graph generation.
We provide a large-scale LLM-generated salient event graph dataset NYT-SEG with three major event relation types for distant supervision ( documents), along with a human-annotated test set ( documents).
@@ 2 Related Work
The early idea of event relation graph construction comes from UzZaman et al. (2013), which introduces a dataset for evaluating an end-to-end system which takes raw text as input and output TimeML annotations (i.e., temporal relations). CAEVO McDowell et al. (2017) and Cogcomptime Ning et al. (2018) both utilise a wide range of manually designed features to train MaxEnt and averaged perception for extracting events and relations. Han et al. (2019b) proposed a joint event and relation extraction model based on BERT Devlin et al. (2019) and BiLSTM Panchendrarajan and Amaresan (2018). Other researchers focus on developing specialised sub-systems to classify extracted event pairs for relations Ning et al. (2019); Han et al. (2019a); Wang et al. (2020); Pergola et al. (2021a, b); Tan et al. (2021, 2023). ATOMIC Sap et al. (2019) is a large-scale commonsense knowledge graph containing the causes and effects of events. MAVEN-ERE Wang et al. (2022) is built with event coreference, temporal, causal and subevent relations. However, ATOMIC and MAVEN-ERE completely rely on crowdsourcing and thus are difficult to extend. MAVEN-ERE is less than half the size of our dataset and does not consider the saliency of events.
@@ 3 Cascading LLMs to Generate Salient Event Graphs
CALLMSAE combines various prompts in a pipelined manner to generate salient event graphs. In this section, we introduce the prompts for generating salient events, and then the method for generating relation graphs based on the salient events. Lastly, we define an evaluation metric for comparing event graphs: Hungarian Graph Similarity.
@@ 3.1 Generate Salient Events
The summarisation test (Section 1) is often used to guide the annotation of salient events or entities Dunietz and Gillick (2014); Liu et al. (2018); Lu et al. (2023); Lyu and Pergola (2024b). These studies identify events or entities included in human-written summaries as salient. Similarly, we instruct LLMs to generate a summary first and then extract events from it. Examples of the summarisation and salient event generation prompts are shown in the Appendix (Table 11 and 12).
@@ 3.2 Generate Graphs as Code Completion
While LLMs can extract salient events, they often struggle with identifying event relations Chan et al. (2023); Sun et al. (2022, 2024); Tan et al. (2024). Prompt engineering for extracting event relations is complex due to the need to incorporate various terminologies and graph constraints. Moreover, prompt efficiency is crucial as generating a large-scale dataset with LLMs can still incur significant computational costs, albeit less than crowdsourcing.
In our method, the main prompt for generating the event relation graph is formulated as a Python code completion task. The graph is defined using the NetworkX333https://networkx.org/documentation/stable/ package in Python, with nodes representing the salient events generated in Section 3.1. LLMs are instructed to complete the code by adding relation edges using NetworkX’s APIs.
Recent research suggests that formulating prompts as code can enhance LLMs’ reasoning abilities Wang et al. (2023); Zhang et al. (2023). In our task, the Python code format effectively incorporates all necessary terminologies, enabling LLMs to understand them without confusion. The Python code format also allows for the inclusion of constraints (e.g., ensuring the graph is a directed acyclic graph) and additional instructions (e.g., ask for explanations) as comments. LLMs can generate explanations as comments without disrupting the main content of the graph, which is difficult to achieve in other formats (e.g., JSON). Moreover, the code template simplifies parsing the response, as LLMs are directed to use the “.add_edge()” function to add the relations.
@@ 3.4 Complement Relation Types
Hierarchical, temporal, and causal relations are not independent of each other. In this case, we found that providing the graph for the first relation can benefit the generation of the dependent relation’s graph. Specifically, we predict the hierarchical relation graph first. Then, we provide this graph to the LLMs and ask them to generate the temporal relation graph. Lastly, with both the hierarchical and temporal relation graphs available, the LLMs predict the causal relation graph.
The hierarchical relation describes two closely related events at different granularity levels. It focuses on the inherent semantics of the events and does not depend on other relation types. For example, “writing a dissertation” is a subevent of “doing a PhD”. Therefore, we choose to predict the hierarchical relations first. Temporal relations can depend on hierarchical relations. For example, knowing “doing a PhD” happened before “being prompted to Professor” allows us to deduce that “writing a thesis” also happened before “being prompted to Professor”. Thus, we predict temporal relations after hierarchical relations. Lastly, causal relations depend on both hierarchical and temporal relations, as the antecedent event in a causal relation must occur before the consequence. Therefore, the causal relation is predicted in the last step. For more details about the entire prompting process, please refer to the pseudocode in Appendix D.
@@ 3.5 Hungarian Graph Similarity
It is challenging to compare event relation graphs generated by LLMs due to the abstractive nature of generation, making it difficult to align the generated events with the gold standard events Li et al. (2023). Moreover, salient events are often high-level and abstract rather than fine-grained and concrete, which means some variations in wording are not only acceptable but also expected. Instead of using exact matching Zhao et al. (2024) or rule-based token matching Tan et al. (2024) on events and relations to calculate , adopting semantic-based evaluation metrics is more reasonable and fair. As more tasks adopt text generation frameworks, many researchers are also turning to metrics based on language models rather than traditional token matching metrics like ROUGE and BLUE Goyal et al. (2022); Pratapa et al. (2023); Lyu and Pergola (2024a).
In this study, we propose a novel metric for evaluating LLM-generated event graphs, called Hungarian Graph Similarity (HGS). The metric is based on the Hungarian assignment algorithm Kuhn (1955), which is widely used in object detection to match generated objects and target objects Carion et al. (2020). It can find the optimal assignment given a cost matrix containing the distance between elements in two lists of objects. We adapt this algorithm to match predicted edges with edges in the gold standard graphs as follows:
@@
Encode the events using SFR-Embedding-Mistral Meng et al. (2024).
@@ 4 Dataset
In this section, we describe how we construct the NYT-SEG which includes LLM-generated distant training pairs and a human-annotated dataset based on the New York Times (NYT) corpus.
@@ 4.2 Annotation Settings
We recruited annotators from Prolific444http://www.prolific.com. There are two subtasks: salient event identification and event relation identification. In the first subtask, the participants are asked to identify the salient event triplets: actor, trigger, and object (optional). We provide the definition of an event and several examples in the guidelines. They are instructed to do the summarisation test: the salient events should be the events they would include in the summary of the given article. Moreover, we provide some prominent features for helping annotators to identify salient events Choubey et al. (2018); Jindal et al. (2020):
@@ 5 Experiments
| Mean event number | Event Frequency | First Appearance | Stretch Size | |
@@
5.1 Model Settings
We compare against the following baselines:
CAEVO McDowell et al. (2017) is a pipeline system based on a MaxEnt classifier and manual features for extracting events and relations.
Madaan and Yang (2021) trained language models on CAEVO-generated linearised graphs with LM objective. We implemented their method to train a Flan-T5 model.
GPT-4 OpenAI, (2024) and GPT-3.5 are based on the generative pre-train framework. We used “gpt-4-1106-preview” and “gpt-3.5-turbo” respectively.
Mixtral is an LLM based on the Mistral model and the mixture of expert framework. We used the 8x7B instruct Jiang et al. (2024).
Llama3 is an LLM based on the Llama framework. We used the 70B-instruct 8bit version for a balance between speed and performance.
We fine-tuned a Flan-T5-base (M) with the relation graphs generated by CALLMSAE, following the same method as in Tan et al. (2024). The baseline prompt evaluates whether each event pair is supported by the document, akin to the hallucination grader described in Section 3.3. Thus, it serves as an ablation of our method without incorporating the code prompt. Another baseline, which asks for the relation type given an event pair, is also tested and included in our experiment.
5.2 Event Saliency Evaluation
@@
| Code Prompt (dependent rels) | N.A. | N.A. | N.A. | ||||||
@@ 5.3 Salient Event Relation Graph Evaluation
The salient event relation graph evaluation results are shown in Table 5. The compared methods (row - ) are outperformed by Baseline Prompt (row ) on all relation types. Baseline Prompt (rel type) (row ), which asks the model to generate all possible relation types given the event pair, performs slightly better than Baseline Prompt. However, Baseline Prompt and Baseline Prompt (relation type) are slow and costly because the number of prompts they need for building one graph is , where is the number of events in the document. On the other hand, the time complexity of Code Prompt is . Moreover, Code Prompt’s overall HGS is significantly higher than Baseline Prompt and Baseline Prompt (relation type) on all relation types. Baseline Prompt check the event pairs more thoroughly and thus have higher recall but its precision is much lower. The complete CALLMSAE combines the code prompt and hallucination grader for iterative refinement, checking missing relations and verifying them to prevent hallucination. Code Prompt (dependent rels) (the th row) is an ablation of CALLMSAE (the row), differing only in the absence of iterative refinement. These results highlight the effectiveness of the hallucination grading approach, which effectively increases the precision and strikes a balance with recall.
@@
| Hier | Temp | Causal | Overall | |
| Baseline (rel type) | ||||
| CALLMSAE | ||||
| Fine-tuned T5 |
5.4 Human Evaluation of Event Graph
We recruited additional annotators to evaluate the generated graph on of the test documents. We asked them whether the relation edges are correct. These scores () can be viewed as precision (Table 6). The human evaluation correlates with the HGS scores, verifying our conclusion.
| Format Error | Cycle | |
| GPT-3.5 | % | % |
| GPT-4 | % | % |
| Mixtral | % | % |
| Llama3 | % | % |
@@ 5.5 Format Error and Cycles in the Graphs
We specified the relation graphs as directed acyclic graphs in the prompt. If there is a cycle in the generated graph, it means that the LLM failed to follow the instructions. A cycle also indicates constraint violations because all the relations in the graphs are asymmetric. We use the APIs in NetworkX package to detect cycles in the transitive closure of the graphs. If the Python interpreter returns an error, it is classified as a format error. We prompt each LLM three times on the test set. Table 7 shows the average number of documents encountering format errors or cycles. All LLMs have low rates of format errors which shows that state-of-the-art LLMs can understand the instruction well and generate executable Python code. Among them, GPT-3.5 and Llama3 have no errors. About of graphs generated by GPT-3.5 have cycles, suggesting that GPT-3.5 may have inferior reasoning ability compared to other LLMs. GPT-4 and Mixtral both have low rates of cycle occurrence, but they are beaten by Llama3 which has no cycle in all generations, showing its remarkable understanding of the transitive and asymmetric constraints in the complex event relation graphs.